Cornerstone Exercise 16.5 (Algorithmic)
Multiple-Product Break-Even and Target Profit

Vandenberg, Inc., produces and sells two products: a ceiling fan and a table fan. Vandenberg plans to sell 40,000 ceiling fans and 80,000 table fans in the coming year. Product price and cost information includes:

Ceiling Fan Table Fan
Price $58   $13  
Unit variable cost $16   $9  
Direct fixed cost $22,400   $47,000  

Common fixed selling and administrative expenses total $86,000.

Required:

1. What is the sales mix estimated for next year (calculated to the lowest whole number for each product)?
Sales mix of ceiling fans to table fans =  :

2. Using the sales mix from Requirement 1, form a package of ceiling fans and table fans. How many ceiling fans and table fans are sold at break-even? Do not round intermediate calculations. If required, round your final answers to the nearest whole number.

Break-even ceiling fans
Break-even table fans


Hide

3. Prepare a contribution-margin-based income statement for Vandenberg, Inc., based on the unit sales calculated in Requirement 2. If an amount is zero, enter "0". Enter any negative product margin and losses with a minus sign. Do not round intermediate calculations. Round your final answers to nearest dollar.

Vandenberg, Inc.

Contribution-Margin-Income Statement

For the Coming Year

Ceiling Fans

Table Fans

Total

$  

$  

$  

  

  

  

$  

$  

$  

  

  

  

$  

$  

$  

  

$  

4. What if Vandenberg, Inc., wanted to earn operating income equal to $14,800? Calculate the number of ceiling fans and table fans that must be sold to earn this level of operating income. (Hint: Remember to form a package of ceiling fans and table fans based on the sales mix and to first calculate the number of packages to earn an operating income of $14,800.) Round your intermediate calculations and final answers to nearest number.

Break-even ceiling fans
Break-even table fans
Answer #1
Sales mix of ceiling fans to table fan 0.5
Ceiling Fan Table Fan
Price per unit 58 13
Variable cost 16 9
Contribution Margin per unit 42 4
Sales Mix 0.333333333 0.666666667
Weighetd average CM 14 2.666666667 16.66666667
Break even fans = Fixed costs/Weighted average CM per unit
                                                                                                 9,324.00 fans
Ceiling Fans              3,108
Table Fans              6,216
Contribution Margin Income statement
Ceiling Fan Table Fan Total
Revenue          180,264            80,808          261,072
Variable costs            49,728            55,944          105,672
Contribution Margin          130,536            24,864          155,400
Direct Fixed cost            22,400            47,000            69,400
Segment Margin          108,136          (22,136)            86,000
Common fixed cost            86,000
Net operating income 0
Total fans required       10,212.00
Ceiling Fans              3,404
Table Fans              6,808

А -40000/80000 1 Sales mix of ceiling fans to table fan 2 3 Ceiling Fan Table Fan =C4-C5 16 =B4-B5 =0.5/1.5 =B6 B7 =1/1.5 =C6